# Towards scalable active steering protocols for genuinely entangled state manifolds

### v1

@@ Abstract
We introduce and analyze an active steering protocol designed to target multipartite entangled states. The protocol involves multiple qubits subjected to weak Bell pair measurements with active feedback, where the feedback operations are optimized to maximize the Quantum Fisher Information. Our scheme efficiently reaches a genuinely entangled one-parameter state manifold. Numerical simulations for systems with up to 20 qubits suggest that the protocol is scalable and allows high multipartite entanglement across the system.
Introduction.—Active steering protocols have recently attracted a lot of attention [1, 2, 3, 4, 5, 6, 7, 8]. By a sequence of (weak) measurements, followed by feedback operations determined by the measurement outcomes, one may prepare, stabilize, or manipulate arbitrary quantum states. (Following Ref. [9], we use “steering” as proxy for “guiding” the system, which differs from “quantum steering” in quantum information theory [10].) Weak, i.e., almost non-invasive, measurements can be performed, for instance, by weakly coupling each system qubit to its own detector qubit, with projective measurements of the detector qubits [11]. The feedback policy is typically based on a cost function. Previous active steering protocols, see, e.g., Refs. [4, 6, 8], have been limited to $N\leq 6$ system qubits, mainly because practically useful fidelity-based cost functions imply an exponential scaling of the algorithmic demands with system size ($N$).
@@
Below, we formulate the active steering protocol in a platform-independent manner. However, since weak measurements play a key role in the protocol, a physical realization with fast qubit readout is desirable, e.g., superconducting Andreev qubits with detector readout times of order 10 ns [22]. As shown in Ref. [8], active steering is robust against error channels with sufficiently weak error rate. For simplicity, we thus neglect external noise and study the idealized error-free case. Our protocol assumes that the system is initialized at time $t=0$ in a simple product state, say, ${|\Psi(t=0)\rangle=|000\cdots\rangle}$, and that for a given measurement record, the state trajectory $|\Psi(t)\rangle$ can be stored and updated on a classical computer for each time step of the quantum protocol. In the conclusions, we briefly discuss how the state tracking requirement may be avoided altogether in modified versions of our scheme.
@@
Figure 1: Schematic time evolution of the active steering protocol. We show three time steps for $N=6$ system qubits (red squares) coupled by steering operators $H_{n}$ (straight vertical lines) to their own detector qubits (grey circles). The qubit chain has periodic boundary conditions. A possible scheme for the Bell measurements of neighboring detector qubit pairs in subsequent cycles is indicated.
Protocol and QFI.—We schematically illustrate the protocol in Fig. 1. We consider $N$ system qubits described by Pauli matrices $\sigma_{n}^{\alpha}$ (with $\alpha=x,y,z$ and $n=1,\ldots,N$), where each system qubit couples only to its own detector qubit described by Pauli matrices $\tau_{n}^{\beta}$. We neither allow for direct couplings between different system qubits nor between different detector qubits, while the Hamiltonian $H_{n}$ (“steering operator”) for the $n$th system-detector qubit pair can be selected from the set of Pauli gates,
@@
with $\alpha_{n}\in\{x,y,z\}$ and $\beta_{n}\in\{x,z\}$. For simplicity, we assume a fixed coupling $J$ and degenerate zero-energy states for all uncoupled qubits. The steering parameters $K_{n}=(\alpha_{n},\beta_{n})$ are chosen according to a decision making scheme in every time step of the protocol as described below, with $[H_{n},H_{n^{\prime}}]=0$ for arbitrary $K_{n}$ and $K_{n^{\prime}}$.
The protocol starts at time $t=0$ by initializing all system and detector qubits in $|0\rangle$ and $|0\rangle_{d}$ (the subscript $d$ refers to detector qubit space), respectively, i.e., the system state is $|\Psi(t=0)\rangle=|000\cdots\rangle$. We then group the $N$ qubits into neighboring pairs ($n,n+1$), see Fig. 1, where all subsequent operations for different pairs commute and can thus be performed simultaneously. (For odd $N$, one “idle” qubit remains whose location is chosen from a uniform random distribution.) Given the state $|\Psi(t)\rangle$, we select the steering couplings for this pair, $(K_{n},K_{n+1})$, such that the measurement-averaged expectation value of the QFI after a time step of duration $\delta t$ is maximized.
@@
For collective observables ${\cal O}=\frac{1}{2}\sum_{n=1}^{N}O_{n}$, where $O_{n}$ are local operators, the QFI can be used to probe the multipartite entanglement structure of the state $|\Psi\rangle$ [13, 14, 15]. If the QFI satisfies the inequality $F_{Q}>mN$, then at least $(m+1)$ parties of the system are entangled. Namely, $m\leq N$ represents the size of the biggest entangled block. The upper bound $F_{Q}\sim N^{2}$ corresponds to the so-called genuinely multipartite entanglement, and it is attained by the states in Eq. (1).
In our protocol, we parametrize $O_{n}={\bf s}_{n}\cdot\bm{\sigma}_{n}$, with ${{\bf s}_{n}=(s_{n}^{x},s_{n}^{y},s_{n}^{z})}$ an arbitrary unit vector and ${\bm{\sigma}_{n}=(\sigma_{n}^{x},\sigma_{n}^{y},\sigma_{n}^{z})}$. Using the optimal choice for $(K_{n},K_{n+1})$, see the discussion after Eq. (Towards scalable active steering protocols for genuinely entangled state manifolds) below, one time-evolves the coupled system-plus-detector system for a time step $\delta t$. Next, a projective measurement of the detector qubit pair is done in its Bell basis $\{|\Phi_{\xi,\eta}\rangle_{d}\}$ [21], where $|\Phi_{\xi=0,\eta=\pm}\rangle_{d}=\left(|00\rangle_{d}\pm|11\rangle_{d}\right)/\sqrt{2}$ and $|\Phi_{\xi=1,\eta=\pm}\rangle_{d}=\left(|01\rangle_{d}\pm|10\rangle_{d}\right)/\sqrt{2}$. Symmetric ($\eta=+1$) and antisymmetric ($\eta=-1$) Bell states have even $(\xi=0)$ or odd ($\xi=1$) parity, where “symmetry” refers to qubit exchange while even (odd) “parity” means that states are built from the basis $\{|00\rangle_{d},|11\rangle_{d}\}$ ($\{|01\rangle_{d},|10\rangle_{d}$}). Such measurements can be implemented by commuting measurements of the Pauli operators $\tau_{n}^{z}\tau_{n+1}^{z}=\pm 1$ and $\tau_{n}^{x}\tau_{n+1}^{x}=\pm 1$. Finally, one re-initializes all detector qubits in the state $|0\rangle_{d}$ and iterates the protocol until convergence has been achieved. Since the initial detector state $|00\rangle_{d}$ (for each time step and each qubit pair) has even parity, measurement outcomes with odd parity ($\xi=1$) are referred to as quantum jumps. The above measurements realize entanglement swapping [23, 24, 25, 26, 27, 28, 29, 30] and tend to increase entanglement in the system state $|\Psi(t)\rangle\to|\Psi(t+\delta t)\rangle$, see Ref. [6] for a detailed discussion.
@@
with $d\rho=|d\Psi\rangle\langle\Psi|+|\Psi\rangle\langle d\Psi|+|d\Psi\rangle\langle d\Psi|$ and ${\rho=|\Psi\rangle\langle\Psi|}$. (Here $\langle A\rangle_{\rm ms}$ denotes a measurement average of the quantity $A$ using the probabilities $p_{\xi,\eta}$.) For practical calculations, Eq. (Towards scalable active steering protocols for genuinely entangled state manifolds) is evaluated by using a Bloch tensor state representation [6]. We compute $\langle dF_{Q}\rangle_{\rm ms}$ for all possible steering parameters and then choose $(K_{n},K_{n+1})$ such that $\langle dF_{Q}\rangle_{\rm ms}$ is maximized. In our numerical simulations, the system state is propagated according to the SSE (4). In contrast to Ref. [6], the time evolution is not terminated once a certain threshold value for $F_{Q}$ has been reached but the quantum state trajectory continues evolving according to the above protocol.

### v2

@@ Abstract
We introduce and analyze an active steering protocol designed to target multipartite entangled states. The protocol involves multiple qubits subjected to weak Bell pair measurements with active feedback, where the feedback operations are optimized to maximize the Quantum Fisher Information. Our scheme efficiently reaches a genuinely entangled one-parameter state manifold. Numerical simulations for systems with up to 22 qubits suggest that the protocol is scalable and allows high multipartite entanglement across the system.
I Introduction
Active steering protocols have recently attracted a lot of attention [1, 2, 3, 4, 5, 6, 7, 8]. By a sequence of (weak) measurements, followed by feedback operations determined by the measurement outcomes, one may prepare, stabilize, or manipulate arbitrary quantum states. (Following Ref. [9], we use “steering” as proxy for “guiding” the system, which differs from “quantum steering” in quantum information theory [10].) Weak, i.e., almost non-invasive, measurements can be performed, for instance, by weakly coupling each system qubit to its own detector qubit, with projective measurements of the detector qubits [11]. The feedback policy is typically based on a cost function. Previous active steering protocols, see, e.g., Refs. [4, 6, 8], have been limited to $N\leq 6$ system qubits, mainly because practically useful fidelity-based cost functions imply an exponential scaling of the algorithmic demands with system size ($N$).
@@ I Introduction
In contrast to the active steering protocol discussed below, which employs minimally invasive weak measurements, recent works have explored measurement-based protocols with a single round of strong (projective) Bell measurements plus feedback, see Refs. [22, 23, 24] and references therein. Such protocols allow one to prepare a broad family of genuinely multipartite entangled states, including the GHZ state (1), in a very efficient manner. Depending on the experimental platform at hand, it is nonetheless of interest to study active steering protocols since one can in principle access arbitrary target states, see Ref. [6] and our discussion below. Moreover, active steering protocols offer the potential for generalization to quantum systems with continuous degrees of freedom.
In Sec. II, we formulate the active steering protocol in a platform-independent manner. Weak measurements and the associated quantum feedback effects have by now become standard experimental tools which are employed in various platforms, see, e.g., Refs. [25, 26, 27, 28, 29, 30, 31, 32, 33]. They allow for high-precision measurements, and an application of these techniques to our protocol could be very promising. Moreover, as shown recently in Ref. [8] for one or two system qubits, by including amplitude damping and dephasing in the stochastic master equation, active steering schemes of the type considered below are expected to be robust against the presence of error channels with sufficiently weak error rate. In fact, Ref. [8] established the existence of a finite threshold error rate, and as long as the error rate stays below the threshold, errors can be corrected “on the fly” by the protocol. Albeit numerical simulations of the present version of our protocol including error channels are computationally prohibitively expensive for large $N$ because one has to simulate the time evolution of mixed states, the fact that the error threshold rate is almost identical for $N=1$ and $N=2$ [8] suggests that a finite error threshold will persist at least for moderate values of $N$. If the system dynamics is described by a mixed state, one has to implement an unraveling procedure in practice for numerical stability. For large $N$, such approaches become exponentially costly in terms of computational demands. This restriction also applies when the detectors have non-ideal measurement efficiency since such effects are also captured by describing the system dynamics in terms of mixed states [34]. Furthermore, the QFI becomes more complicated for mixed states. For simplicity, we thus neglect external noise channels and non-ideal measurement efficiencies, and study the error-free case with ideal measurements below. However, a modified implementation of our protocol that may allow to circumvent these restrictions is discussed in Sec. IV.
Since weak measurements play a key role in our protocol, a physical realization with fast qubit readout is desirable, e.g., superconducting Andreev qubits with detector readout times of order 10 ns [35] could offer a good option [36]. We note that active steering protocols of similar type have recently been experimentally implemented [37, 38]. Our protocol assumes that the system is initialized at time $t=0$ in a simple product state, say, ${|\Psi(t=0)\rangle=|000\cdots\rangle}$, and that for a given measurement record, the state trajectory $|\Psi(t)\rangle$ can be stored and updated on a classical computer for each time step of the quantum protocol. After introducing the protocol in Sec. II, we present numerical simulation results in Sec. III. We primarily focus on the GHZ state, but we also show how to prepare so-called Dicke states using this approach. In Sec. IV, we discuss several open issues and directions for future research. In particular, we outline how the state tracking requirement may be avoided in modified implementations of our protocol.
II Protocol and QFI
@@
Figure 1: Schematic time evolution of the active steering protocol. (a) We show three time steps for $N=6$ system qubits (red squares) coupled by steering operators $H_{n}$ (straight vertical lines) to their own detector qubits (grey circles). The qubit chain has periodic boundary conditions. A possible scheme for the Bell measurements of neighboring detector qubit pairs in subsequent cycles is indicated. (b) We mainly consider the case of Bell measurements of nearest-neighbor detector qubit pairs. (c) In Fig. 5 below, we also study the case where arbitrary detector qubit pairs can be subjected to Bell measurements (full connectivity).
choose $|\Psi(0)\rangle=|0\dots 0\rangle$ and $\mathcal{O}$
Classical computer Quantum computer $|\Psi(t)\rangle$ $|\Psi(t)\rangle\rightarrow|\Psi(t)\rangle\otimes|00\rangle_{d}$downarrow $\downarrow$ $\mathrm{max}_{K\in\mathcal{K}}\langle dF_{Q}(|\Psi\rangle,K)\rangle_{\rm ms}$ $\overset{K\in\mathcal{K}}{\longrightarrow}$ $e^{-i\delta tH_{K}}|\Psi(t)\rangle\otimes|00\rangle_{d}$downarrow $\downarrow$ $|\Psi(t+\delta t|\xi,\eta)\rangle$ ${(\xi,\eta)}$ ${}_{d}\langle\Phi_{\xi,\eta}|e^{-i\delta tH_{K}}|00\rangle_{d}\times|\Psi(t)\rangle$
Table 1: Active steering protocol using the QFI as cost function. A classical computation of the measurement-averaged cost function change $\langle dF_{Q}(|\Psi\rangle,K)\rangle_{\rm ms}$ determines the optimal coupling $K=(K_{n},K_{n+1})$ out of the coupling family $\mathcal{K}=\{K_{n}=(\alpha_{n},\beta_{n})|\alpha_{n}\in\{x,y,z\};\beta_{n}\in\{x,z\}\}$ for qubit pair $(n,n+1)$. This choice is fed into the quantum computer as system-detector coupling. After unitary time evolution of duration $\delta t$, the detector is measured in the Bell basis $|\Phi_{\xi,\eta}\rangle_{d}$, where $\xi$ and $\eta$ correspond to the possible measurement outcomes. These are fed back into the classical computer to update and keep track of the system state.
We schematically illustrate the protocol in Fig. 1 and Table 1. Each protocol step of time duration $\delta t$ has two components, namely (i) unitary evolution of the coupled system and detector qubits under the chosen feedback Hamiltonian, followed by (ii) weak measurements of the system qubits via Bell pair measurements of the detector qubits. Depending on the measurement outcomes, the feedback Hamiltonian for the next iteration step is then determined according to the decision making scheme discussed below.
We consider $N$ system qubits described by Pauli matrices $\sigma_{n}^{\alpha}$ (with $\alpha=x,y,z$ and $n=1,\ldots,N$), where each system qubit couples only to its own detector qubit described by Pauli matrices $\tau_{n}^{\beta}$. We neither allow for direct couplings between different system qubits nor between different detector qubits, while the Hamiltonian $H_{n}$ (“steering operator”) for the $n$th system-detector qubit pair can be selected from the set of Pauli gates,
@@
with $\alpha_{n}\in\{x,y,z\}$ and $\beta_{n}\in\{x,z\}$. For simplicity, we assume a fixed coupling $J$, where we put $J=+1$ in what follows, and degenerate zero-energy states for all uncoupled qubits. The steering parameters $K_{n}=(\alpha_{n},\beta_{n})$ are chosen according to a decision making scheme in every time step of the protocol as described below, with $[H_{n},H_{n^{\prime}}]=0$ for arbitrary $K_{n}$ and $K_{n^{\prime}}$.
Table 1 illustrates the active steering protocol using Bell state measurements on neighboring pairs, see Fig. 1(b), of detector qubits. Later on, we also consider the case in Fig. 1(c) where Bell measurements can be performed for arbitrarily chosen pair orderings (where a given pairing order is determined from a uniform random distribution), but we focus on the nearest-neighbor case in what follows. For a given time step of the protocol, all non-overlapping pairs $(n,n+1)$ can be steered simultaneously, see Fig. 1(a). One can either choose an alternating sequence of pairings on subsequent time steps, as shown in Fig. 1(a), or simply assign the pairing order randomly. In our simulations, we found the second option to be more efficient. As explained below, the active decision making is performed on a classical computer using the knowledge about the present state of the system.
Let us now describe the protocol in detail. We start at time $t=0$ by initializing all system and detector qubits in $|0\rangle$ and $|0\rangle_{d}$ (the subscript $d$ refers to detector qubit space), respectively, i.e., the system state is $|\Psi(t=0)\rangle=|000\cdots\rangle$. We then group the $N$ qubits into neighboring pairs ($n,n+1$), see Fig. 1, where all subsequent operations for different pairs commute and can thus be performed simultaneously. (For odd $N$, one “idle” qubit remains whose location is chosen from a uniform random distribution.) Given the state $|\Psi(t)\rangle$, for active steering towards the GHZ state, we select the steering couplings for this pair, $(K_{n},K_{n+1})$, such that the measurement-averaged expectation value of the QFI after a time step of duration $\delta t$ is maximized.
@@
For collective observables ${\cal O}=\frac{1}{2}\sum_{n=1}^{N}O_{n}$, where $O_{n}$ are local operators, the QFI can be used to probe the multipartite entanglement structure of the state $|\Psi\rangle$ [13, 14, 15]. If the QFI satisfies the inequality $F_{Q}>mN$, then at least $(m+1)$ parties of the system are entangled. Namely, $m\leq N$ represents the size of the biggest entangled block. The upper bound $F_{Q}\sim N^{2}$ corresponds to the so-called genuinely multipartite entanglement. In particular, the family of states which saturate the maximum value of the multipartite entanglement is an arbitrary superposition of the eigenvectors of ${\cal O}$ with largest and smallest eigenvalues. For ${\cal O}=\frac{1}{2}\sum_{n=1}^{N}\sigma^{z}_{n}$, these correspond to the states defined in Eq. (1).
In our protocol, we parametrize $O_{n}={\bf s}_{n}\cdot\bm{\sigma}_{n}$, with ${{\bf s}_{n}=(s_{n}^{x},s_{n}^{y},s_{n}^{z})}$ an arbitrary unit vector and ${\bm{\sigma}_{n}=(\sigma_{n}^{x},\sigma_{n}^{y},\sigma_{n}^{z})}$. Using the optimal choice for $(K_{n},K_{n+1})$, discussed below after Eq. (II), one time-evolves the coupled system-plus-detector system for a time step $\delta t$. Next, a projective measurement of the detector qubit pair is done in its Bell basis $\{|\Phi_{\xi,\eta}\rangle_{d}\}$ [21], where $|\Phi_{\xi=0,\eta=\pm}\rangle_{d}=\left(|00\rangle_{d}\pm|11\rangle_{d}\right)/\sqrt{2}$ and $|\Phi_{\xi=1,\eta=\pm}\rangle_{d}=\left(|01\rangle_{d}\pm|10\rangle_{d}\right)/\sqrt{2}$. Symmetric ($\eta=+1$) and antisymmetric ($\eta=-1$) Bell states have even $(\xi=0)$ or odd ($\xi=1$) parity, where “symmetry” refers to qubit exchange while even (odd) “parity” means that states are built from the basis $\{|00\rangle_{d},|11\rangle_{d}\}$ ($\{|01\rangle_{d},|10\rangle_{d}$}). Such measurements can be implemented by commuting measurements of the Pauli operators $\tau_{n}^{z}\tau_{n+1}^{z}=\pm 1$ and $\tau_{n}^{x}\tau_{n+1}^{x}=\pm 1$. Finally, one re-initializes all detector qubits in the state $|0\rangle_{d}$ and iterates the protocol until convergence has been achieved. Since the initial detector state $|00\rangle_{d}$ (for each time step and each qubit pair) has even parity, measurement outcomes with odd parity ($\xi=1$) are referred to as quantum jumps. The above measurements realize entanglement swapping [39, 40, 41, 42, 43, 44, 45, 46] and tend to increase entanglement in the system state $|\Psi(t)\rangle\to|\Psi(t+\delta t)\rangle$, see Ref. [6] for a detailed discussion.
@@
with $d\rho=|d\Psi\rangle\langle\Psi|+|\Psi\rangle\langle d\Psi|+|d\Psi\rangle\langle d\Psi|$ and ${\rho=|\Psi\rangle\langle\Psi|}$. (Here $\langle A\rangle_{\rm ms}$ denotes a measurement average of the quantity $A$ using the probabilities $p_{\xi,\eta}$.) We compute $\langle dF_{Q}\rangle_{\rm ms}$ for all possible steering parameters and then choose $(K_{n},K_{n+1})$ such that $\langle dF_{Q}\rangle_{\rm ms}$ is maximized. We emphasize that $(K_{n},K_{n+1})$ is determined separately for each qubit pair at a given time step. In our numerical simulations, the system state is propagated according to the SSE (4). In contrast to Ref. [6], the time evolution is not terminated once a certain threshold value for $F_{Q}$ has been reached but the quantum state trajectory continues evolving according to the above protocol.
Let us provide some details on the calculation of $\langle dF_{Q}\rangle_{\rm ms}$. The classical computation of $\langle dF_{Q}(|\Psi\rangle,K)\rangle_{\rm ms}$ for $H_{K}=H_{n}+H_{n+1}$ with $K=(K_{n},K_{n+1})$ is performed numerically in a Bloch tensor representation of the system state,
$|\Psi\rangle\langle\Psi|=\frac{1}{2^{N}}\sum_{\cal S}R_{\cal S}{\cal S},\quad R_{\cal S}=\langle\Psi|{\cal S}|\Psi\rangle,$ (7)
with the Pauli string operator ${\cal S}=\sigma_{1}^{\mu_{1}}\sigma_{2}^{\mu_{2}}\cdots\sigma_{N}^{\mu_{N}}$ with $\mu_{j}\in\{0,1,2,3\}$. Using this representation to parametrize the observables, the average QFI change is obtained as
$\langle dF_{Q}\rangle_{\rm ms}$ $=\sum_{n\neq m}\sum_{\alpha_{i},\alpha_{j}}s_{n}^{\alpha_{n}}s_{m}^{\alpha_{m}}\langle dQ_{n,m}^{\alpha_{n},\alpha_{m}}\rangle_{\rm ms}-2\left(\sum_{n,\alpha_{n}}s_{n}^{\alpha_{n}}\langle dR_{n}^{\alpha_{n}}\rangle_{\rm ms}\right)\left(\sum_{n,\alpha_{n}}s_{n}^{\alpha_{n}}R_{n}^{\alpha_{n}}\right)-2\delta t\sum_{\eta}\frac{1}{\langle c_{\eta}^{\dagger}c_{\eta}\rangle}\left(\sum_{n,\alpha_{n}}s_{n}^{\alpha_{n}}G^{(\eta)}_{\alpha_{n}}\right)^{2},$ (8)
where we used the reduced single-qubit Bloch vectors $R_{n}^{\alpha_{n}}=\langle\Psi|\sigma_{n}^{\alpha_{n}}|\Psi\rangle$ and the two-qubit correlators $Q_{n,m}^{\alpha_{n},\alpha_{m}}=\langle\Psi|\sigma_{n}^{\alpha_{n}}\sigma_{m}^{\alpha_{m}}|\Psi\rangle$. For explicit expressions for $\langle dR_{n}^{\alpha_{n}}\rangle_{\rm ms}$ and $\langle dQ_{n,m}^{\alpha_{n},\alpha_{m}}\rangle_{\rm ms}$, see Eq. (27) in Ref. [6]. The second-order state change of the single qubit density matrices for $\xi=1$ measurement outcomes is given by
$G^{(\eta)}_{\mu_{i}}$ $=-\sum_{m=n,n+1}\sum_{\alpha\neq\alpha_{i}}\Gamma_{m}\delta_{\beta_{m},x}\delta_{\mu_{i},\alpha}R_{\mu_{i}}+\eta\sqrt{\Gamma_{n}\Gamma_{n+1}}\delta_{\beta_{n},x}\delta_{\beta_{n+1},x}\left(\mathcal{F}_{\mu_{i}}-Q_{n,n+1}R_{\mu_{i}}\right),$ (9)
$\mathcal{F}_{\cal S}$ $=\frac{1}{2^{N+1}}\sum_{{\cal S}^{\prime}}R_{{\cal S}^{\prime}}{\rm Tr}\left((\sigma_{n}^{\alpha_{n}}{\cal S}^{\prime}\sigma_{n+1}^{\alpha_{n+1}}+\sigma_{n+1}^{\alpha_{n+1}}{\cal S}^{\prime}\sigma_{n}^{\alpha_{n}}){\cal S}\right).$
With a chosen coupling $(K_{n},K_{n+1})$ for the following time step, the quantum system is evolved accordingly, and the detector is measured subsequently. The stochastic outcome of this measurement is recorded and used to keep track of the running state on the classical computer.
We note that for the preparation of non-GHZ state families, one may choose a specific target value of the QFI, $F_{Q}^{*}$, and then minimize the cost function $|F_{Q}-F_{Q}^{\ast}|$. This approach allows one to use active steering protocols targeting different classes of state manifolds associated with a specific QFI value. As an example, we show results for Dicke states in Sec. III below.
III Numerical simulation results